# ShapeLLM: Universal 3D Object Understanding for Embodied Interaction

### v2

@@ 5 Discussions
5.1 What is learned from multi-view alignment?
Fig. 6 illustrates the visualization of the attention map in the last cross-attention layer, documenting the image query to which each local patch in the attention map primarily attends. It provides evidence that multi-view alignment achieves geometrically informed spatial understanding, which may implicitly encompass the estimation of the object pose and a more profound knowledge of 3D spatial relationships.
Table 5: 3D referring expression grounding on GAPartNet. Accuracy with an IoU threshold of 0.25 is reported. ${}^{\dagger}$: Fine-tuned on GAPartNet images. ${}^{\ddagger}$: Inference with 3 in-context demonstrations.
@@ 5.1 What is learned from multi-view alignment?
LLaVA-13B [93] 1-View Img. 0.0 0.0 0.0 0.0 0.0 0.0 0.0
LLaVA-13B [93] 4-View Img. 0.0 0.0 0.0 0.0 0.0 0.0 0.0
${}^{\dagger}$ [93] 1-View Img. 1.8 9.3 3.8 0.0 2.1 11.1 4.4
${}^{\dagger}$ [93] 4-View Img. 2.5 13.7 7.7 0.0 4.3 11.1 6.2
GPT-4V [115] 4-View Img. 0.0 0.0 0.0 0.0 0.0 0.0 0.0
${}^{\ddagger}$ [115] 4-View Img. 0.1 1.6 0.0 0.0 0.0 0.0 0.3
ShapeLLM-7B Point Cloud 5.9 25.8 11.5 3.4 5.1 11.1 10.5
ShapeLLM-13B Point Cloud 7.6 26.7 11.5 6.7 6.8 11.1 11.7
5.2 Is ShapeLLM grounded in physical worlds?
Tab. 5 compares ShapeLLM with image-only methods on 3D referring expression grounding (REG) of 6-DoF poses on GAPartNet. The results show that: i) Image-only methods cannot perform zero-shot geometry-necessary 6-DoF pose estimation. ii) Compared to image-only methods with 2D to 6-DoF pose estimation fine-tuning or in-context prompting, ShapeLLM still performs significantly better. It demonstrates the necessity of geometry and the difficulty of the ill-posed 2D to 6-DoF pose estimation problem.
@@ 5.3 Can ShapeLLM generalize to unseen objects?
Fig. 7 shows the part understanding examples of unseen objects. While ShapeLLM’s 6-DoF pose estimation is trained on GAPartNet, which primarily consists of indoor articulated furniture. It has demonstrated promising generalization potential of spatial understanding on the open-world objects, paving ways for scaling up spatial-awareness training.
@@ 6 Related Works
Interaction-oriented 3D Understanding  Interaction with 3D objects typically involves concept-only interaction and physical-grounded interaction [15]. The former works focus on 3D perception and semantic parsing, such as 3D object recognition and scene perception [126, 127, 157, 159, 99]. By utilizing language for open-ended interaction in 3D, a number of works demonstrate successful 3D scene QA [174, 109], grounding [16], and captioning [17]. Recently, some works propose to utilize foundation models like LLMs or CLIP for open-ended 3D object recognition [186, 198, 94] and scene segmentation [181, 122]. Guo & Zhang et al. [54] utilizes ImageBind [49] and LLaMA-Adapter [187] to realize point cloud-based interactive QA. Following LLaVA, PointLLM [166] conducts supervised fine-tuning by constructing a visual instruction-following dataset. Other works focus on scene-level tasks utilizing comprehensive 2D features [199, 68] or 3D features distilled from 2D images into LLMs [62, 199, 68]. The second kind of interaction typically requires physical understanding in 3D, such as part understanding [105, 113, 47, 96], 6-DoF pose estimation [177, 156, 84, 97, 160], particularly useful for human-object interaction (HOI) and robotic manipulation [95, 128, 167, 154, 46, 178, 50, 45, 48, 139, 85, 47] and complex robotic planning [39, 34, 72, 90, 14, 141]. In this work, we focus on both physical and conceptual interactions with 3D shapes for embodied understanding.
Multimodal Large Language Models  Multimodal comprehension, which allows human interaction with textual and visual elements, has witnessed significant advancements, particularly in extending LLMs like LLaMA [150, 149, 21]. The early efforts predominantly revolved around integrating LLMs with various downstream systems by employing it as an agent [161, 57, 172, 88, 140, 171, 148, 6, 155]. Significant success has been demonstrated within this plugin-style framework. Due to the remarkable capabilities of LLMs, aligning the visual semantic space with language through parameter-efficient tuning [64, 2, 83, 187, 197, 173] and instruction tuning [168, 93, 24, 37] has emerged as the prevailing approach in current research. To further enhance interactive capabilities, some approaches have been developed towards visual-interactive multimodal comprehension by precisely referring to instruction tuning [191, 123, 19, 189]. Another family advances the developments of LLMs endowed with content creation beyond comprehension, notable efforts include DreamLLM [37], GILL [81], Emu [147, 145], SEED [44], NeXt-GPT [162], and Kosmos-G [118].
Figure 7: Part understanding examples of unseen objects.
@@ 7 Conclusions
This paper presents ShapeLLM, a 3D multimodal LLM for embodied interaction, capable of generalizable recognition and embodied interaction comprehension. We first propose a novel 3D point cloud encoder, ReCon++, by utilizing multi-view distillation and scaling up 3D representation learning, which serves as the foundation 3D representation encoder for ShapeLLM. Then, we perform 3D visual instruction tuning on constructed instruction-following data for general and embodied comprehension. We also established a 3D evaluation benchmark, 3D MM-Vet, severing as assessing the 4-level capacity in embodied interaction scenarios, varying from basic perception to control statements generation.
Appendix A Additional Experiments
A.1 Ablation Study
A.1.1 ShapeLLM Architecture
Architecture
@@ A.1.1 ShapeLLM Architecture
The local and transformation-invariant 3D geometric embeddings $\mathbf{x}_{i}=\mathop{\text{MAX}}\limits_{\mathbf{p}_{i,j}\in{\mathcal{N}_{i}}}\big{(}\Phi_{\gamma}\left(\xi_{i,j}\right)\big{)}$ for $\mathcal{P}_{i},i=1,2,\dots,N_{s}$ is used as 3D token embeddings of ReCon++, where $\Phi_{\gamma}$ is a per-point MLP point feature extractor [126, 127] and $\xi_{i,j}$ is the feature of $j$-th neighbour point $\mathbf{p}_{i,j}$ in the neighbourhood $\mathcal{N}_{i}$. Let $\{\mathbf{g}_{q}^{\text{image}}\}_{q=1}^{G}$ be $G$ multi-view image global queries and $\mathbf{g}^{\text{text}}$ be the global text query. ReCon++ outputs the local and global 3D point cloud representations by taking 3D embeddings and global queries as inputs:
@@
Visual Prompt Number  Fig. 8 shows the performance of ShapeLLM using different numbers of prompts, including 1, 8, 16, 32, and 64. This ablation study has shown that a different number of prompts leads to varied improvements, and the optimal setting is 32. This observation is similar to VPT [76] where the prompts used to modulate Transformer attention should be studied [59].
@@
A.1.2 Baseline Improvement
Can we improve the baseline to bridge the gap between PointLLM and ShapeLLM? In Tab. 6, we study two technical factors that are contributed by ShapeLLM: 3D point cloud encoder and SFT data.
  • Improvement from encoder. (Line 1) First, by changing PointLLM’s encoder to ReCon++, a significant improvement of +4.20% is obtained. This demonstrates the significantly better 3D representation extraction of ReCon++ compared to ULIP-2. It is consistent with previous findings in Tab. 1 and Tab. 2 that ReCon++ outperforms ULIP-2 by a large margin regarding 3D representation transferring learning and zero-shot learning.
  • Improvement from data. (Line 2) As stated in Sec. 2.2, we have constructed instruction-following data for supervised fine-tuning (SFT) using GPT-4V involving comprehensive topics. By further using the SFT data curated by us, PointLLM’s performance gap to ShapeLLM has been fulfilled. This demonstrates the superiority of our SFT data, where the decent quality comes from the more advanced GPT4-V model using multi-view images and the comprehensive topics covered in the data.
Table 7: Ablation study on baseline improvements. Results are tested on 3D MM-Vet with the baseline model PointLLM-13B [166] using different point cloud encoders and SFT data. Line 0 is the result with the original PointLLM configuration. Line 1 and 2 denote the results of changing the point cloud encoder to ReCon++ and using our SFT data. Note that when ReCon++ is used as the encoder, $\mathbf{E}_{\text{APE}}$ and $\mathbf{E}_{\text{prompt}}$ are also used by default.
Encoder SFT Data Rec Know Gen Spat Emb Total
0 ULIP-2 [170] PointLLM 46.6 48.3 38.8 45.2 50.9 46.6
1 ReCon++ PointLLM 47.5 52.8 43.6 44.9 54.5 50.8
2 ReCon++ Ours 46.8 53.0 53.9 45.3 68.4 53.1
A.2 Multimodal Comprehension with ShapeLLM
@@
InstructBLIP-7B [24] 1-View Img. 25.51 43.50 45.34 47.41 48.48
InstructBLIP-13B [24] 1-View Img. 28.69 34.25 44.97 45.90 48.86
LLaVA-7B [93] 1-View Img. 39.71 50.00 46.71 45.61 47.10
LLaVA-13B [93] 1-View Img. 36.59 51.75 38.28 46.37 45.90
3D-LLM [62] 3D Obj. + Mul.-V. Img. - 45.25 33.42 44.48 43.68
PointLLM-7B [166] Point Cloud 52.63 53.00 44.85 47.47 48.55
PointLLM-13B [166] Point Cloud 52.78 54.00 48.15 47.91 49.12
ShapeLLM-7B Point Cloud 53.08 54.50 46.92 48.20 49.23
ShapeLLM-13B Point Cloud 52.96 54.00 48.94 48.52 49.98
Figure 9: 3D multimodal dialogue using single-view point cloud inputs. All answers are generated by ShapeLLM-13B with single-view occluded inputs. ShapeLLM achieves outstanding robustness against such occlusion, which is the most commonly met in real worlds.
Singe-View Point Cloud Inputs  As stated in Sec. 4.2, we construct 3D MM-Vet-C which studies three kinds of corruptions commonly met in real-world scenarios: “single-view”, “jitter”, and “rotate”. Among these corruptions, the “singe-view” issue stands out as the most critical challenge since obtaining the objects’ complete point clouds is non-trivial, similar to multi-view images. As a result, everyday real-world robots only get single-view 3D perceptions with sensors such as RGB-D [56]. Fig. 9 shows the qualitative examples of ShapeLLM-13B’s response using single-view point cloud inputs, demonstrating surprisingly outstanding robustness in processing such occluded inputs.
Human Win Rate Comparison  GPT-4 [116] is widely used as an evaluator in natural language and vision language processing, as seen in recent modern benchmarks like MM-Bench and MM-Vet. Recent studies [163] have demonstrated that ChatGPT-based evaluation is more closely aligned with human preferences compared to traditional metrics. With GPT4-turbo, the standard deviation of 3D MM-Vet is less than 0.1. To further verify the soundness of the models’ response, we also conduct human evaluation and report the win rate in Fig. 10, where ShapeLLM demonstrates superior preference by humans.
@@
Visual Hallucination
@@
Figure 11: Qualitative comparison examples of visual hallucination.
A.3 Representation Transferring with ReCon++
Linear SVM Evaluation
Linear SVM evaluation [8, 152] can be used to evaluate the discriminative quality of pretrained features [51]. The results on ModelNet40 are shown in Tab. 9. It shows that our ReCon++ outperforms Point-BERT, which also uses plain Transformers with contrastive objectives, by a clear margin of +6.2%. Compared to hierarchical Transformers methods, our ReCon++ outperforms PointM2AE [185] by +0.7%.
@@ A.3 Representation Transferring with ReCon++
Few-Shot 3D Object Recognition
Few-shot learning is critical for evaluating the representation transferring capabilities in data and training efficiency. We conduct few-shot 3D object recognition experiments on the ModelNet40 [164] dataset, and the results are shown in Tab. 10. Our ReCon++ achieves state-of-the-art performance in all the benchmarks compared to previous works.
@@
Appendix B Additional Information about 3D MM-vet
B.1 Evaluation System
@@
B.2 Analysis
The 3D MM-Vet evaluation benchmark consists of 5 different categories of questions. In Fig. 12 we report the distribution of problem categories. The knowledge and General Visual Recognition parts contain multiple subparts that comprehensively evaluate these capacities and thus hold higher proportions. Fig. 13 shows an example of how we prompt GPT-4 for 3D MM-Vet evaluation. Fig. 14 and Fig. 15 illustrate additional examples of 3D MM-Vet Q&As.
Figure 14: Additional Visualization example of 3D MM-Vet Q&A pairs.
Figure 15: Additional Visualization example of 3D MM-Vet Q&A pairs.
Appendix C Implementation details
ReCon++  Following the standard ViT [38] architecture, we design four different model structures consistent with prior work [129, 94, 195]. The model parameters are shown in Tab. 11. Following OpenShape [94], we employ four datasets as pretraining data, namely Objaverse [29], ShapeNet [13], ABO [22], and 3D-FUTURE [42]. Each point cloud sample has a size of 10000$\times$6, where the first three dimensions represent $xyz$ coordinates, and the latter three dimensions represent $rgb$ values.
Table 11: Details of ReCon++ model variants. This table format follows Dosovitskiy et al..

### v3

@@ 5 Discussions
5.1 Is ShapeLLM grounded in physical worlds?
Tab. 6 compares ShapeLLM with image-only methods on 3D referring expression grounding (REG) of 6-DoF poses on GAPartNet [50]. The results show that: i) Image-only methods cannot perform zero-shot geometry-necessary 6-DoF pose estimation. ii) Compared to image-only methods with 2D to 6-DoF pose estimation fine-tuning or in-context prompting, ShapeLLM still performs significantly better. It demonstrates the necessity of geometry and the difficulty of the ill-posed 2D to 6-DoF pose estimation problem, as well as the importance of using 3D point clouds as input for spatial intelligence.
Table 6: 3D referring expression grounding on GAPartNet [50]. Accuracy with an IoU threshold of 0.25 is reported. : Fine-tuned on GAPartNet images. : Inference with 3 in-context demonstrations.
@@ 5.1 Is ShapeLLM grounded in physical worlds?
LLaVA-13B [96] 1-View 2D Image 0.0 0.0 0.0 0.0 0.0 0.0 0.0
LLaVA-13B [96] 4-View 2D Image 0.0 0.0 0.0 0.0 0.0 0.0 0.0
LLaVA-13B [96] 1-View 2D Image 1.8 9.3 3.8 0.0 2.1 11.1 4.4
LLaVA-13B [96] 4-View 2D Image 2.5 13.7 7.7 0.0 4.3 11.1 6.2
GPT-4V [120] 4-View 2D Image 0.0 0.0 0.0 0.0 0.0 0.0 0.0
GPT-4V [120] 4-View 2D Image 0.1 1.6 0.0 0.0 0.0 0.0 0.3
ShapeLLM-7B 3D Point Cloud 5.9 25.8 11.5 3.4 5.1 11.1 10.5
ShapeLLM-13B 3D Point Cloud 7.6 26.7 11.5 6.7 6.8 11.1 11.7
@@ 5.2 Can ShapeLLM generalize to unseen objects?
Figure 7: Part understanding examples of unseen objects beyond GAPartNet.
Fig. 7 shows the part understanding examples of unseen objects. While ShapeLLM’s 6-DoF pose estimation and spatial awareness are trained on GAPartNet, which primarily consists of indoor articulated furniture objects. It has demonstrated promising generalization potential of spatial understanding on the open-world objects, paving ways for scaling up spatial awareness training.
@@ 6 Related Works
Interaction-Oriented 3D Understanding  Interaction with 3D objects typically involves concept-only interaction and physical-grounded interaction [15]. The former works focus on 3D perception and semantic parsing, such as 3D object recognition and scene perception [132, 133, 163, 165, 104]. By utilizing language for open-ended interaction in 3D, a number of works demonstrate successful 3D scene QA [180, 111], grounding [16], and captioning [17]. Recently, some works propose to utilize foundation models like LLMs or CLIP for open-ended 3D object recognition [193, 206, 97, 37] and scene segmentation [188, 127]. Guo & Zhang et al. [57] utilizes ImageBind [52] and LLaMA-Adapter [194] to realize point cloud-based interactive QA. Following LLaVA, PointLLM [172] conducts supervised fine-tuning by constructing a visual instruction-following dataset. Other works focus on scene-level tasks utilizing comprehensive 2D features [207, 71] or 3D features distilled from 2D images into LLMs [65, 207, 71]. The second kind of interaction typically requires physical understanding in 3D, such as part understanding [108, 116, 50, 98], 6-DoF pose estimation [181, 162, 88, 100, 166], particularly useful for human-object interaction (HOI) and robotic manipulation [99, 134, 173, 160, 49, 184, 53, 48, 51, 145, 89, 50, 117, 20] and complex robotic planning [40, 35, 74, 93, 14, 147]. In this work, we focus on both physical and conceptual interactions with 3D shapes for embodied understanding.
Multimodal Large Language Models  Multimodal comprehension, which allows human interaction with textual and visual elements, has witnessed significant advancements, particularly in extending LLMs like LLaMA [156, 155, 22]. The early efforts predominantly revolved around integrating LLMs with various downstream systems by employing it as an agent [167, 60, 178, 91, 146, 177, 154, 6, 161]. Significant success has been demonstrated within this plugin-style framework. Due to the remarkable capabilities of LLMs, aligning the visual semantic space with language through parameter-efficient tuning [67, 2, 86, 194, 205, 179] and instruction tuning [174, 96, 25, 36] has emerged as the prevailing approach in current research. To further enhance interactive capabilities, some approaches have been developed towards visual-interactive multimodal comprehension by precisely referring to instruction tuning [199, 129, 19, 196]. Another family advances the developments of LLMs endowed with content creation beyond comprehension [36, 84, 151, 153, 47, 168, 123].
@@ 7 Conclusions
This paper introduces ShapeLLM, the first 3D MLLM for embodied interaction, excelling in generalizable recognition and interaction comprehension. We present ReCon++, a novel 3D point cloud encoder leveraging multi-view distillation and advanced 3D representation learning, forming the basis for ShapeLLM. We perform 3D visual instruction tuning on curated instruction-following data for broad and embodied comprehension. Additionally, we establish 3D MM-Vet, a benchmark to evaluate four levels of capacity in embodied interaction scenarios, from fundamental recognition to control statement generation.
Acknowledgments
The work was supported by the Dushi Program from Tsinghua University, the National Key R&D Program of China (2022YFB2804103), and the National Science and Technology Major Project of China (2023ZD0121300).
Appendix 0.A Additional Experiments
0.A.1 ShapeLLM Architecture
@@
The local and transformation-invariant 3D embeddings $\mathbf{x}_{i}=\mathop{\text{MAX}}\limits_{\mathbf{p}_{i,j}\in{\mathcal{N}_{i}}}\big{(}\Phi_{\gamma}\left(\xi_{i,j}\right)\big{)}$ for $\mathcal{P}_{i}^{s},i=1,2,\dots,N_{s}$ is used as 3D token embeddings of ReCon++, where $\Phi_{\gamma}$ is a per-point MLP point feature extractor [132, 133] and $\xi_{i,j}$ is the feature of $j$-th neighbour point $\mathbf{p}_{i,j}$ in the neighbourhood $\mathcal{N}_{i}$. Let $\{\mathbf{g}_{q}^{\text{image}}\}_{q=1}^{G}$ be $G$ multi-view image global queries and $\mathbf{g}^{\text{text}}$ be the global text query. ReCon++ outputs the local and global 3D point cloud representations by taking 3D embeddings and global queries as inputs:
@@ 0.A.1.1 Input Components
0.A.1.2 Visual Prompt Number
Fig. 8 shows the performance of ShapeLLM using different numbers of prompts, including 1, 8, 16, 32, and 64. This ablation study has shown that a different number of prompts leads to varied improvements, and the optimal setting is 32. This observation is similar to VPT [79] where the prompts used to modulate Transformer attention should be studied [62].
@@ 0.A.1.2 Visual Prompt Number
0.A.2 Multimodal Comprehension with ShapeLLM
@@
InstructBLIP-7B [25] 1-View 2D Image 25.51 43.50 45.34 47.41 48.48
InstructBLIP-13B [25] 1-View 2D Image 28.69 34.25 44.97 45.90 48.86
LLaVA-7B [96] 1-View 2D Image 39.71 50.00 46.71 45.61 47.10
LLaVA-13B [96] 1-View 2D Image 36.59 51.75 38.28 46.37 45.90
3D-LLM [65] 3D Object + Multi-View 2D Image - 45.25 33.42 44.48 43.68
PointLLM-7B [172] 3D Point Cloud 52.63 53.00 44.85 47.47 48.55
PointLLM-13B [172] 3D Point Cloud 52.78 54.00 48.15 47.91 49.12
ShapeLLM-7B 3D Point Cloud 53.08 54.50 46.92 48.20 49.23
ShapeLLM-13B 3D Point Cloud 52.96 54.00 48.94 48.52 49.98
0.A.2.1 Human Win Rate Comparison
GPT-4 [119] is widely used as an evaluator in natural language and vision language processing, as seen in recent modern benchmarks like MM-Bench and MM-Vet. Recent studies [169] have demonstrated that ChatGPT-based evaluation is more closely aligned with human preferences compared to traditional metrics. With GPT4-turbo, the standard deviation of 3D MM-Vet is less than 0.1. To further verify the soundness of the models’ response, we also conduct human evaluation and report the win rate in Fig. 9, where ShapeLLM demonstrates superior preference by humans.
@@ 0.A.2.1 Human Win Rate Comparison
Figure 10: Qualitative comparison examples of visual hallucination.
0.A.2.2 Visual Hallucination
@@
Table 9: Language-only baseline results on 3D MM-Vet.
Method Rec Know Gen Spat Emb Total
LLaMA2-7B-Chat 11.8 10.6 22.1 14.6 25.8 16.2
GPT-3.5-Turbo 2.9 5.4 17.3 10.2 27.8 11.7
GPT-4-Turbo 1.7 3.6 16.1 6.6 26.0 9.8
ShapeLLM-7B 36.5 35.1 36.9 35.1 52.5 39.0
ShapeLLM-13B 46.8 53.0 53.9 45.3 68.4 53.1
0.A.2.3 Language-Only Baseline
To verify whether the capabilities of our model stem solely from large language models rather than from 3D perception abilities, we also conducted a Language-Only Baseline evaluation. This involved testing the performance of 3D MMVet on large language models such as ChatGPT [118] and LLaMA [156]. Results is shown in Tab. 9. Our intentional usage of pronouns (e.g., “it”) for referring to unknown objects in questions leads to poor performance, and LLaMA hallucinates but scores higher than ChatGPT, which rejects to answer without specific inputs Fig. 11.
Figure 11: Language-only baseline analysis.
0.A.3 Representation Learning with ReCon++
Linear SVM  Linear SVM evaluation [8, 158] can be used to evaluate the discriminative quality of pretrained features [54]. The results on ModelNet40 are shown in Tab. 10. The results show that our ReCon++ outperforms both plain and hierachical Transformer methods by a clear margin.
@@ 0.A.3 Representation Learning with ReCon++
Few-Shot 3D Object Recognition  Few-shot learning is critical for evaluating the representation transferring capabilities in data and training efficiency. We conduct few-shot 3D object recognition experiments on ModelNet40, and the results are shown in Tab. 11. Our ReCon++ achieves state-of-the-art performance in all the benchmarks compared to previous works.
@@
0.A.3.1 Multi-view Alignment visualization analysis.
Fig. 12 illustrates the visualization of the attention maps in the last cross-attention layer, documenting the image query to which each local patch primarily attends. It provides evidence that multi-view alignment achieves geometrically informed spatial understanding, which may implicitly encompass the estimation of the object pose and a more profound knowledge of 3D spatial relationships.
0.A.3.2 ReCon++ Key Modifications Analysis
We conduct an ablation study on the two key modifications of ReCon++, namely scaling up and multi-view alignment, and the results are presented in Tab. 12. The results demonstrate that: i) scaling up 3D representation is critical for both 3D representation learning, and stonger 3D representation understanding brought by ReCon++ consistently yields better 3D multimodal comprehension; ii) the proposed multi-view distillation further leads to significant improvement.
Table 12: Ablation study on scaling and multi-view alignment.
scaling multi-view Zero-Shot 3D MM-Vet
6.7 15.8
10.3 21.9
51.5 48.2
53.7 53.1
Appendix 0.B Additional Information about 3D MM-vet
0.B.1 Evaluation System
@@
0.B.2 Analysis
The 3D MM-Vet evaluation benchmark consists of 5 different categories of questions. In Fig. 13 we report the distribution of problem categories.
@@ 0.B.2 Analysis
The knowledge and General Visual Recognition parts contain multiple subparts that comprehensively evaluate these capacities and thus hold higher proportions. Fig. 14 shows an example of how we prompt GPT-4 for 3D MM-Vet evaluation. Fig. 15 and Fig. 16 illustrate additional examples of 3D MM-Vet Q&As.
Table 13: Sample categories of 3D MM-Vet data.
Category Characters Life Art Architecture Animals
Number 11 16 10 13 9
0.B.3 ChatGPT Costs
In constructing the Supervised Finetuning dataset for ShapeLLM and conducting inference on 3D MMVet using the GPT-4 or GPT-4V API, we have roughly estimated the costs. For ShapeLLM’s training data, which contains over 50k Q&A pairs, with each request yielding 5-6 Q&A pairs, the estimated cost is approximately \$900. As for inference on 3D MMVet, with only 232 samples and averaging five requests per sample, the cost is estimated to be around \$12.
Table 14: ReCon++ model variants, which follow ViT [39].